# Homogeneous Projective Coordinates for the Bondi-Metzner-Sachs Group

### v1

@@
ABSTRACT This paper studies the Bondi-Metzner-Sachs group in homogeneous projective coordinates, because it is then possible to write all transformations of such a group in a manifestly linear way. The 2-sphere metric, Bondi-Metzner-Sachs metric, asymptotic Killing vectors, generators of supertranslations, as well as boosts and rotations of Minkowski spacetime, are all re-expressed in homogeneous projective coordinates. Last, the integral curves of vector fields which generate supertranslations are evaluated in detail. This work prepares the ground for more advanced applications of the differential geometry of asymptotically flat spacetimes in projective coordinates.
@@ 1 Introduction
The Bondi-Metzner-Sachs [1, 2, 3] asymptotic symmetry group of asymptotically flat spacetime has received again much attention over the last decade by virtue of its relevance for black-hole physics [4, 5, 6], the group-theoretical structure of general relativity [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] and the infrared structure of fundamental interactions [21, 22, 23, 24]. The appropriate geometric framework can be summarized as follows. In spacetime models for which null infinity can be defined, the cuts of null infinity are spacelike two-surfaces orthogonal to the generators of null infinity [25]. On using the familiar stereographic coordinate
@@
$\left(\begin{matrix}1&0&a\cr 0&1&b\cr 0&0&1\end{matrix}\right)\left(\begin{matrix}x\cr y\cr z\end{matrix}\right)=\left(\begin{matrix}x+a\cr y+b\cr 1\end{matrix}\right),$ (1.14)
@@
The work in Ref. [19] has outlined the resulting geometric picture, where $(w_{0},w_{1},w_{2})$ are viewed as homogeneous coordinates in a complex projective plane. In our paper we have instead a less abstract and more concrete task: since the Bondi-Metzner-Sachs transformation (1.2) becomes linear when expressed in terms of $z_{0}$ and $z_{1}$, we are aiming to develop the Bondi-Metzner-Sachs formalism with the associated Killing vector fields by using the pair of variables $(z_{0},z_{1})$ instead of $(\zeta,{\overline{\zeta}})$. For this purpose, the homogeneous projective coordinates for the 2-sphere are studied in Sect. 2, while the Bondi-Sachs metric in homogeneous coordinates is considered in Sect. 3. Asymptotic Killing fields for supertranslations are evaluated in Sect. 4, while their flow is investigated in Sect. 5. Concluding remarks and open problems are presented in Sect. 6, while technical details are provided in the Appendices.
@@ 2 Homogeneous coordinates on the $2$-sphere
$\tan(\theta/2)==\dfrac{2\sin(\theta/2)\cos(\theta/2)}{2\cos^{2}(\theta/2)}=\dfrac{\sin(\theta)}{1+\sqrt{1-\sin^{2}(\theta)}},$ (2.3)
@@
$d\theta^{2}$ $=$ $\dfrac{d\theta}{dz_{0}}\dfrac{d\theta}{dz_{0}}dz_{0}^{2}+\dfrac{d\theta}{dz_{1}}\dfrac{d\theta}{dz_{1}}dz_{1}^{2}+2\dfrac{d\theta}{dz_{0}}\dfrac{d\theta}{dz_{1}}dz_{0}dz_{1}$ (2.4)
@@
$\sin^{2}(\theta)\,d\varphi^{2}$ $=$ $4z^{2}_{0}z^{2}_{1}\,d\varphi^{2}=4z^{2}_{0}z^{2}_{1}\left\{\dfrac{d\varphi}{dz_{0}}\dfrac{d\varphi}{dz_{0}}dz_{0}^{2}+\dfrac{d\varphi}{dz_{1}}\dfrac{d\varphi}{dz_{1}}dz_{1}^{2}+2\dfrac{d\varphi}{dz_{0}}\dfrac{d\varphi}{dz_{1}}dz_{0}dz_{1}\right\}$ (2.5)
@@
-- $4z^{2}_{1}\left(\frac{1-4z^{2}_{0}z^{2}_{1}+2\sqrt{1-4z^{2}_{0}z^{2}_{1}}}{1-4z^{2}_{0}z^{2}_{1}}\right)dz_{0}^{2}+8z_{0}z_{1}dz_{0}dz_{1}$
@@ 3 Bondi-Sachs metric in homogeneous coordinates
On passing from $(\theta,\varphi)$ to $(z_{0},z_{1})$ coordinates, we find the metric components of (3.1) expressed as follows:
@@
The Bondi gauge $\partial_{r}\det\left(r^{-2}g_{AB}\right)=0$ implies that [26] $\gamma^{AB}C_{AB}=0$, where $\gamma^{AB}$ is given in Eq. (2.9). With our coordinates, this relation reads as
@@
Now we are ready to evaluate the BMS generators in homogeneous coordinates in order to determine the supertranslations.
@@ 4 Asymptotic Killing fields
Moreover, the preservation of the Bondi gauge condition yields [26]
@@
From these relations one can calculate the four components of $\xi^{\mu}$. At this stage, instead of repeating the detailed calculations already available, for example, in Ref. [26], we can compute the asymptotic Killing fields in homogeneous coordinates by using the familiar transformation law of vector fields. In other words, the work in Ref. [26] has defined the stereographic variable (we write $\psi$ rather than $z$ used in Ref. [26], in order to avoid confusion with our $\zeta$ in Eq. (1.1))
@@
and has found, in Bondi coordinates $u,r,\theta,\varphi$, the asymptotic Killing fields $\xi_{T}^{+}$ where the components depend on a function $f$ and on the Bondi coordinates. On denoting as usual by $Y_{l}^{m}$ the spherical harmonics on the $2$-sphere, one finds [26]
@@ 5 Flow of supertranslation vector fields
In order to appreciate that the familiar geometric constructions are feasible also in projective coordinates, we now consider the flow of supertranslation vector fields (4.9)-(4.11). For example, by virtue of (2.7), and defining $p=(u,r,z_{0},z_{1})$, the task of finding the flow of the supertranslation vector fields (4.9), (4.10) and (4.11) consists of solving a system of nonlinear and coupled differential equations. For this purpose, we denote by $\sigma,\Sigma,\chi$, respectively, the appropriate flow, and define
@@ 6 Concluding remarks and open problems
As far as we can see, the interest of our investigation lies in having shown that homogeneous projective coordinates lead to a fully computational scheme for all applications of the BMS group. This might pay off when more advanced properties will be studied. In particular, we have in mind the concept of superrotations [21, 22, 26] on the one hand, and the physical applications of the Segre manifold advocated in Ref. [19]. In other words, since our Eq. (1.15) contains Eq. (1.12), which in turn is just a re-expression of the BMS transformation (1.2), one might aim at embedding the study of BMS symmetries into the richer mathematical framework of complex analysis [27] and algebraic geometry. Our paper has tried to prepare the ground for such a synthesis.
@@ Acknowledgments
The boost ($K_{i}$) and rotation ($J_{ij}$) vector fields for Lorentz transformations in Minkowski spacetime can be written in $u,r,\xi,{\bar{\xi}}$ coordinates as is shown, for example, in Refs. [21, 26]. At that stage, by using again Eqs. (4.7), (4.8) and (A7)-(A10) we find
@@
with the help of two additional sets of $14$ nonvanishing functions, one set for each Lie bracket in (B20). For example, in the Lie bracket among $\xi_{2}$ and $\xi_{3}$, the coefficient of ${\partial\over\partial z_{0}}$ is a function

### v2

@@
ABSTRACT This paper studies the Bondi-Metzner-Sachs group in homogeneous projective coordinates, because it is then possible to write all transformations of such a group in a manifestly linear way. The 2-sphere metric, Bondi-Metzner-Sachs metric, asymptotic Killing vectors, generators of supertranslations, as well as boosts and rotations of Minkowski spacetime, are all re-expressed in homogeneous projective coordinates. Last, the integral curves of vector fields which generate supertranslations are evaluated in detail. This work prepares the ground for more advanced applications of the geometry of asymptotically flat spacetimes in projective coordinates, by virtue of the tools provided from complex analysis in several variables and projective geometry.
@@ 1 Introduction
The Bondi-Metzner-Sachs [1, 26, 27] asymptotic symmetry group of asymptotically flat spacetime has received again much attention over the last decade by virtue of its relevance for black-hole physics [4, 5, 6], the group-theoretical structure of general relativity [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] and the infrared structure of fundamental interactions [21, 22, 23, 24, 25]. Moreover, since asymptotic symmetries can provide key constraints on the celestial dual to quantum gravity in flat spacetimes, much work has been devoted to the celestial holography program and related issues [26, 27, 28, 29, 30, 31, 32].
The appropriate geometric framework can be summarized as follows. In spacetime models for which null infinity can be defined, the cuts of null infinity are spacelike two-surfaces orthogonal to the generators of null infinity [33]. On using the familiar stereographic coordinate
@@
$\left(\begin{matrix}1&0&a\cr 0&1&b\cr 0&0&1\end{matrix}\right)\left(\begin{matrix}x\cr y\cr 1\end{matrix}\right)=\left(\begin{matrix}x+a\cr y+b\cr 1\end{matrix}\right),$ (1.14)
@@
Within this extended framework, one can consider two complex projective planes [19]. Let $P$ be a point of the first plane with coordinates $(w_{0},w_{1},w_{2})$, and let $P^{\prime}$ be a point of the second plane, with coordinates $(u_{0},u_{1},u_{2})$. One can now consider the nine products between a complex coordinate of $P$ and a complex coordinate of $P^{\prime}$, i.e.
$Z_{hk}=w_{h}u_{k},\;h,k=0,1,2.$ (1.17)
This equation provides the coordinate description of the Segre manifold [34, 35], which is the projective image of the product of projective spaces. It is a natural tool for accommodating the transformations that reduce to the BMS transformations upon restriction to the unit circle $\Gamma$. It contains a complex double infinity of planes, two arrays of planes, and a complex fourfold infinity of quadrics [19, 34], but its differential geometry is still largely unexplored, as far as we know.
Unlike Ref. [19], we have a more concrete task: since the Bondi-Metzner-Sachs transformation (1.2) becomes linear when expressed in terms of $z_{0}$ and $z_{1}$, we are aiming to develop the Bondi-Metzner-Sachs formalism with the associated Killing vector fields by using the pair of variables $(z_{0},z_{1})$ instead of $(\zeta,{\overline{\zeta}})$. For this purpose, the homogeneous projective coordinates for the 2-sphere are studied in Sect. 2, while the Bondi-Sachs metric in homogeneous coordinates is considered in Sect. 3. Asymptotic Killing fields for supertranslations are evaluated in Sect. 4, while their flow is investigated in Sect. 5. Concluding remarks and open problems are presented in Sect. 6, while technical details are provided in the Appendices.
@@ 2 Homogeneous coordinates on the $2$-sphere
$\tan(\theta/2)=\dfrac{2\sin(\theta/2)\cos(\theta/2)}{2\cos^{2}(\theta/2)}=\dfrac{\sin(\theta)}{1+\sqrt{1-\sin^{2}(\theta)}},$ (2.3)
@@
$d\theta^{2}$ $=$ $\left({\partial\theta\over\partial z_{0}}\right)^{2}dz_{0}^{2}+\left({\partial\theta\over\partial z_{1}}\right)^{2}dz_{1}^{2}+2{\partial\theta\over\partial z_{0}}{\partial\theta\over\partial z_{1}}dz_{0}dz_{1}$ (2.4)
@@
$\sin^{2}(\theta)\,d\varphi^{2}$ $=$ $4z^{2}_{0}z^{2}_{1}\,d\varphi^{2}=4z^{2}_{0}z^{2}_{1}\left\{\left({\partial\varphi\over\partial z_{0}}\right)^{2}dz_{0}^{2}+\left({\partial\varphi\over\partial z_{1}}\right)^{2}dz_{1}^{2}+2{\partial\varphi\over\partial z_{0}}{\partial\varphi\over\partial z_{1}}dz_{0}dz_{1}\right\}$ (2.5)
@@
$=$ $-4z^{2}_{1}\left(\frac{1-4z^{2}_{0}z^{2}_{1}+2\sqrt{1-4z^{2}_{0}z^{2}_{1}}}{1-4z^{2}_{0}z^{2}_{1}}\right)dz_{0}^{2}+8z_{0}z_{1}dz_{0}dz_{1}$
@@ 3 Bondi-Sachs metric in homogeneous coordinates
On passing from $(\theta,\varphi)$ to $(z_{0},z_{1})$ coordinates, we find the metric components of (3.1) expressed as follows (the material from our Eq. (3.2) to our Eq. (3.17) can be obtained from Eqs. (4.33), (4.35) and (4.37) in Ref. [36], which relies in turn upon the work in Ref. [37]):
@@
The Bondi gauge $\partial_{r}\det\left(r^{-2}g_{AB}\right)=0$ implies that $\gamma^{AB}C_{AB}=0$, where $\gamma^{AB}$ is given in Eq. (2.9). With our coordinates, this relation reads as
@@
For the discussion of Bondi’s news tensor, mass and angular momentum aspects we refer again to the work in Refs. [36, 37]. Now we are ready to evaluate the BMS generators in homogeneous coordinates in order to determine the supertranslations.
@@ 4 Asymptotic Killing fields
Moreover, the preservation of the Bondi gauge condition yields
@@
From these relations one can calculate the four components of $\xi^{\mu}$. At this stage, we can compute the asymptotic Killing fields in homogeneous coordinates by using the familiar transformation law of vector fields. In other words, the work in Ref. [22] has defined the stereographic variable (we write $\psi$ rather than $z$ used in Ref. [22], in order to avoid confusion with our $\zeta$ in Eq. (1.1))
@@
and has found, in Bondi coordinates $u,r,\theta,\varphi$, the asymptotic Killing fields $\xi_{T}^{+}$ where the components depend on a function $f$ and on the Bondi coordinates. On denoting as usual by $Y_{l}^{m}$ the spherical harmonics on the $2$-sphere, one finds
@@ 5 Flow of supertranslation vector fields
The analysis in this section does not have a direct impact on unsolved problems, but (as far as we can see) can help the general reader. More precisely, in order to appreciate that the familiar geometric constructions are feasible also in projective coordinates, we now consider the flow of supertranslation vector fields (4.9)-(4.11). For example, by virtue of (2.7), and defining $p=(u,r,z_{0},z_{1})$, the task of finding the flow of the supertranslation vector fields (4.9), (4.10) and (4.11) consists of solving a system of nonlinear and coupled differential equations. For this purpose, we denote by $\sigma,\Sigma,\chi$, respectively, the appropriate flow, and define
@@ 6 Concluding remarks and open problems
As far as we can see, the interest of our investigation lies in having shown that homogeneous projective coordinates lead to a fully computational scheme for all applications of the BMS group. This might pay off when more advanced properties will be studied. In particular, we have in mind the concept of superrotations [22, 23] on the one hand, and the physical applications of the Segre manifold advocated in our Introduction and in Ref. [19]. In other words, since our Eq. (1.15) contains Eq. (1.12), which in turn is just a re-expression of the BMS transformation (1.2), one might aim at embedding the study of BMS symmetries into the richer mathematical framework of complex analysis in several variables [38] and algebraic geometry. The exploitment of the complex analysis approach to algebraic geometry appears promising because the singular points of functions of several complex variables form a continuum (see definitions and theorems in Refs. [34, 38]). The potentialities of this framework for studying e.g. superrotations were unforeseen so far, and deserve careful consideration in our opinion. Our paper has tried to prepare the ground for such a synthesis, even though our calculations are not cumbersome.
Moreover, we would like to mention that the research in Refs. [36, 39, 40] has exploited the fact that one can actually work with a completely arbitrary metric on the asymptotic 2-sphere. By doing so, one can write the on-shell expression of $U,U^{A}$ and $\beta$ in our Sect. 3 in terms of this arbitrary 2-sphere metric. This might therefore provide a way to recover our results when taking the particular case in which the 2-sphere metric is expressed in homogeneous projective coordinates. We are grateful to M. Geiller for this remark, and also for having brought to our attention the work in Ref. [41], where the authors have written the solution space and the asymptotic Killing vectors and their action in the case of an even more general gauge than Bondi-Sachs.
Author Contributions
The authors have equally contributed to conceptual and technical parts of the paper. All authors have read and agreed to the published version of the manuscript.
Funding
This research received no external funding.
Data Availability Statement
No new data were created.
@@ Acknowledgments
Conflicts of Interest
The authors declare no conflict of interest.
@@
The boost ($K_{i}$) and rotation ($J_{ij}$) vector fields for Lorentz transformations in Minkowski spacetime can be written in $u,r,\xi,{\bar{\xi}}$ coordinates as is shown, for example, in Ref. [22]. At that stage, by using again Eqs. (4.7), (4.8) and (A7)-(A10) we find
@@
with the help of two additional sets of $14$ nonvanishing functions, one set for each Lie bracket in (B20). For example, in the Lie bracket among $\xi_{2}$ and $\xi_{3}$, the coefficient of ${\partial\over\partial z_{0}}$ is the function